OpenStudy (anonymous):

What is the equation of the line that passes through (0,-3), (-2,-13) and (2,3

OpenStudy (anonymous):

Assume the equation is of the form y = ax^2 + bx + c. It's not a straight line that goes through those 3 points. Plug in the first pair (0,-3) and you get -3 = a*0^2 +b*0 + c or c = - 3 So now we have y = ax^2 + bx -3 Now plug in the other 2 points we know. You'll get 2 equations in 2 unknowns and you can use the elimination method to solve